Consider the function y = 1.5(2^x)
Make a table for this function, letting x = 2, 1, 0, -1, and -2.

Answer by richwmiller(17219)

just plug in the values 2,1,0,-1,-2 one at a time and calculate y
2^(-1)=1/2
2^(0)=1